i have to find which interval must contain a zero of f with the intermediate value theorem. if i have x^4-x^2-15 how do i get a= and b= and how do i solve that? it s badly described in the book...

Answer by stanbon(75887)

i have to find which interval must contain a zero of f with the intermediate value theorem. if i have x^4-x^2-15 how do i get a= and b= and how do i solve that? it s badly described in the book...
-------------
The idea is common sense. If the function is negative and then becomes
positive it must pass thru zero.
-------------
So, if f(0) = 0 - 9 - 15 = -15
and f(3) = 3^4 - 3^2 - 15 = 81 -24 = 57,
There must be a zero between x = 0 and x = 3
=================================================
Cheers,
Stan H.